OpenStudy (anonymous):

What is the solution to the equation 364c = (8c - 6) ?

OpenStudy (anonymous):

Bring the pronumerals to one side 364c - 8c = -6 356c = -6 c = -6/356 simplify and you will get c = -3/178

OpenStudy (anonymous):

i guess george is right :)

OpenStudy (anonymous):

OOOOHHHH NOOOOO hahaha im sooo sorry its 36^4c= sqrt6^(8c - 6) what happened there?!

OpenStudy (anonymous):

draw it  ?